# Note on shifted primes with large prime factors

### v1

@@ Abstract.
For any $0<c<1$ let
$T_{c}(x)=|\big\{p\leq x:p\in\mathbb{P},P^{+}(p-1)\geq p^{c}\big\}|,$
where $\mathbb{P}$ is the set of primes and $P^{+}(n)$ denotes the largest prime factor of $n$. Erdős proved in 1935 that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\rightarrow 0,\quad\text{as~}c\rightarrow 1,$
where $\pi(x)$ denotes the number of primes not exceeding $x$. Recently, Ding gave a quantitative form of Erdős’ result and showed that for $8/9<c<1$ we have
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq 8\big(c^{-1}-1\big).$
In this article, Ding’s bound is improved to
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c$
@@ 1. Introduction
$T_{c}(x):=\big|\big\{p\leq x:p\in\mathbb{P},P^{+}(p-1)\geq p^{c}\big\}\big|,$
where $0<c<1$ and $\mathbb{P}$ is the set of primes. As an application of the Bombieri-Vinogradov theorem as well as the Brun-Titchmarsh inequality, Goldfeld [13] proved in 1969 that
$\liminf_{x\rightarrow\infty}T_{1/2}(x)/\pi(x)\geq 1/2.$
Goldfeld also pointed out that his arguments could also lead to, for any $c<7/12$
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)>0.$ (1.1)
There are a number of improvements on the value of $c$ in (1.1), see, e.g. Motohashi [23], Hooley [20, 21], Deshouillers–Iwaniec [7], Fouvry [12] and Baker-Harman [2]. The best record of $c$ up to now is $0.677$, obtained by Baker and Harman [3].
In 2015 Luca, Menares and Pizarro-Madariaga [19] considered the explicit lower bound of $T_{c}(x)$ for small values of $c$. Specifically, for $1/4\leq c\leq 1/2$ they proved that
$T_{c}(x)\geq(1-c)\frac{x}{\log x}+E(x)$ (1.2)
where
$E(x)\ll\begin{cases}x\log\log x/(\log x)^{2},&\text{for~}1/4<c\leq 1/2,\\ x/(\log x)^{5/3},&\text{for~}c=1/4.\end{cases}$
Later, Chen and Chen [6] extended the range of $c$ to $(0,1/2)$ in (1.2) with slightly better $E(x)$. Chen and Chen also proved that for any $k\geq 2$ there exists at most one $c\in\big[\frac{1}{k+1},\frac{1}{k}\big)$ such that
$T_{c}(x)=(1-c)\frac{x}{\log x}+o\Big(\frac{x}{\log x}\Big).$
Based on their result, Chen and Chen conjectured that for any $k\geq 1$ and $c\in\big[\frac{1}{k+1},\frac{1}{k}\big)$ we have
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-\frac{1}{k+1}.$ (1.3)
In 2018, Feng and Wu [11] proved that
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-4\int_{1/c-1}^{1/c}\frac{\rho(t)}{t}dt$
holds for $0<c<0.3517\ldots$, where $\rho(u)$ is the Dickman function, defined as the unique continuous solution of the equation differential-difference
$\begin{cases}\rho(u)=1,&0\leq u\leq 1,\\ u\rho^{\prime}(u)=-\rho(u-1),&u>1.\end{cases}$
As a corollary, Feng and Wu proved conjecture (1.3) for $k\geq 3$ by numerical values involving the Dickman function. The lower bounds of $T_{c}(x)$ were further improved by Liu, Wu and Xi [18] to
$\liminf_{x\rightarrow\infty}T_{c}(x)/\pi(x)\geq 1-4\rho(1/c)$
provided $0<c<0.3734\ldots$.
Later Ding [8, Final remarks] pointed out that conjecture (1.3) of Chen and Chen in fact contradicts the Elliott–Halberstam conjecture according to the works of Pomerance [24], Granville [14], Wang [25] and Wu [27]. That is, one has
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)=\lim_{x\rightarrow\infty}T_{c}(x)/\pi(x)=\left(1-\rho\left(\frac{1}{c}\right)\right)\rightarrow 0,\quad\text{as}~c\rightarrow 1$
under the assumption of the Elliott–Halberstam conjecture. Motivated by this, Ding [8] then proved unconditionally that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$ (1.4)
for some absolute constant $c<1$, thus disproving conjecture (1.3) for the case $k=1$. The proof of (1.4) by Ding is based on the following corollary of Brun-Titchmarsh type. For $(a,m)=1$ let $\pi(x;m,a)$ denotes the number of primes $p$ not exceeding $x$ such that $p\equiv a\pmod{m}$.
@@
[27, Lemma 2.2] There exist two functions $K_{2}(\theta)>K_{1}(\theta)>0$, defined on the interval $(0,17/32)$ such that for each fixed $A>0$, and sufficiently large $Q=x^{\theta}$, the inequalities
@@
The constant $c$ in (1.4) could further be specified by explicit values of $K_{1}(\theta)$ in Proposition 1.1. In fact, one has $K_{1}(\theta)\geq 0.16$ for $1/2\leq\theta\leq 13/25$ [2, Theorem 1] and $K_{1}(\theta)\geq 1/100$ for $13/25\leq\theta\leq 17/32$ [22, Eq. (4)]. Using the method of Ding [8] as well as the explicit values of $K_{1}(\theta)$, Xinyue Zang (private communication) obtained that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq\frac{0.496875}{c}<\frac{1}{2},\quad\text{for~}0.993375<c<1.$ (1.5)
However, there are earlier results related to conjecture (1.3) of Chen and Chen as well as Ding’s result (1.4). Actually, as indicated by the proof of a former result of Erdős [10, from line -6, page 212 to line 4, page 213], as early as 1935, people could already conclude that
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\rightarrow 0,\quad\text{as}~c\rightarrow 1$ (1.6)
by combining with Wu’s lemma (see Lemma 2.3 below). Clearly, (1.4) is now a simple corollary of (1.6). In a later article, Ding [9] obtained a quantitative form of Erdős’ result (1.6), stating
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq 8\big(c^{-1}-1\big)$ (1.7)
@@
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<\frac{1}{2}$ (1.8)
@@
[15, page 172, Theorem 5.7] Let $g$ be a natural number, and let $a_{i},b_{i}~(i=1,2,\cdot\cdot\cdot,g)$ be integers satisfying
@@
In this article, we shall give a further improvement of (1.7) with two new ingredients: the first is the employment of Rosser-Iwaniec linear sieve to the prime variable sequence, instead of integer variable polynomial combining with the two dimensional sieve (i.e. Proposition 1.2 above); the second one is that when dealing with the error term coming from linear sieve, we apply a theorem of Bombieri-Friedlander-Iwaniec type with level of distribution $x^{4/7-\varepsilon}$ instead of the classical level $x^{1/2-\varepsilon}$.
@@
$\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)\leq-\frac{7}{2}\log c.$
@@
The following corollary of Theorem 1.1 improved (1.4), (1.5) and (1.8) considerably.
@@
The numerical value of $e^{-\frac{1}{7}}$ is $0.86687\ldots$. In [9, Remarks], it was concluded that $\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$ for any $c>e^{-\frac{1}{2}}=0.60653\ldots$ under the Elliott-Halberstam conjecture. Corollary 1.1 makes some further progress toward this direction.
@@ 2. Fundamental lemmas
Let $\mu(n)$ be the Möbius function. Let $\mathcal{A}$ be a finite sequence of positive integers and $\mathcal{P}$ a subset of primes. For any $z\geq 2$, let
@@
Next for square-free number $d$ with $d|P(z)$, we define
$\mathcal{A}_{d}=:\big\{a\in\mathcal{A}:d|a\big\}.$
Define the sieve function $S(\mathcal{A},\mathcal{P},z)$ to be
$S(\mathcal{A},\mathcal{P},z):=\big|\big\{a\in\mathcal{A}:(a,P(z))=1\big\}\big|.$
Suppose that $\big|\mathcal{A}_{d}\big|$ possesses the following form
where $X$ is an approximation to $|\mathcal{A}|$ and $\omega(d)$ is a multiplicative function satisfying
$0<\omega(p)<p,\quad p\in\mathcal{P}.$ (2.1)
Here, $\omega(d)d^{-1}X$ can be viewed as an approximation of the quantity $\big|\mathcal{A}_{d}\big|$ and $r\big(\mathcal{A},d\big)$ is regarded as the oscillation between $\big|\mathcal{A}_{d}\big|$ and $\omega(d)d^{-1}X$. We also let
$V(z)=:\prod_{\begin{subarray}{c}p|P(z)\end{subarray}}\left(1-\frac{\omega(p)}{p}\right).$
The first lemma is a result of Iwaniec [16, 17] on linear sieve with well factorable error terms. An arithmetic function $\lambda(q)$ is called well factorable of level $Q$ if for any $Q=Q_{1}Q_{2}$, $Q_{1},Q_{2}\geq 1$, there exist two functions $\lambda_{1}$ and $\lambda_{2}$ supported in $1,Q_{1}]$ and $1,Q_{2}]$ respectively such that
@@
Lemma 2.1.
Suppose that there is a constant $K\geq 2$ such that
$\prod_{\begin{subarray}{c}u\leq p<v\\ p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}\leq\frac{\log v}{\log u}\left(1+\frac{K}{\log u}\right)$
for any $v>u\geq 2$. Then for any $\varepsilon>0$ and $D^{1/2}\geq z\geq 2$ we have
$S(\mathcal{A},\mathcal{P},z)\leq XV(z)\left(F\left(\frac{\log D}{\log z}\right)+E\right)+\sum_{h<\exp(8/\varepsilon^{2})}\sum_{d|P(z)}\lambda^{+}_{h}(d)r\big(\mathcal{A},d\big),$
where $sF(s)=2e^{\gamma}$ $(0<s\leq 3)$, $\gamma$ is the Euler constant, and the first error term $E$ satisfies
$E=O\big(\varepsilon+\varepsilon^{-8}e^{K}(\log D)^{-1/3}\big).$
The coefficients $\lambda^{+}_{h}(d)$ satisfy $\big|\lambda_{h}^{+}(d)\big|\leq 1$ and vanish for $d>D$ or $\mu(d)=0$. Especially, $\lambda^{+}_{h}(d)$ are well factorable of level $D$.
Let $\mathbb{N}$ be the set of natural numbers. For $q\in\mathbb{N}$ and $(a,q)=1$, define
$\pi(y;\ell,a,q)=\sum_{\begin{subarray}{c}\ell p\leq y\\ \ell p\equiv a\!\!\!\!\!\pmod{q}\end{subarray}}1,$
where the symbols $p$ will always be primes. As usual, let $\varphi(n)$ be the Euler totient function and ${\rm li}(y)=\int_{2}^{y}\frac{1}{\log t}dt$ be the Gauss function. The second lemma is the following theorem of Bombieri-Friedlander-Iwaniec type given by Wang [26, Proposition 3.2].
@@
$\sum_{\begin{subarray}{c}q\leq Q\\ (a,q)=1\end{subarray}}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leq\ell\leq L_{2}\\ (\ell,q)=1,~2|\ell\end{subarray}}\bigg(\pi(x;\ell,a,q)-\frac{{\rm li}\big(x/\ell\big)}{\varphi(q)}\bigg)\ll\frac{x}{(\log x)^{A}}$
holds for $Q\leq x^{4/7-\varepsilon}$ and $1\leq L_{1}\leq L_{2}\leq x^{1-\varepsilon}$, where the implied constants depend only on $a,A$ and $\varepsilon$.
@@
In fact, the original statement of Wang’s proposition [26, Proposition 3.2] is slightly different from Lemma 2.2. For our applications, we add the additional restriction $2|\ell$ here in Lemma 2.2. The proof is almost the same as Wang [26, Proposition 3.2]. After applying Heath-Brown’s identity, we shall consider the sum
@@
$\sum_{(a,\,q)=1}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leqslant\ell\leqslant L_{2}\\ (\ell,\,q)=1\end{subarray}}f(\ell)\bigg(\pi(x;\,\ell,a,q)-\frac{{\rm{li}}(x/\ell)}{\varphi(q)}\bigg)\ll_{a,A,\varepsilon}\frac{x}{(\log x)^{A}}$
@@
The last lemma is another conjecture of Chen and Chen [6] which was later confirmed by Wu [27, Theorem 2].
@@
$T^{\prime}_{c}(x)=\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ P^{+}(p^{\prime}-1)\geq x^{c}\end{subarray}}1=\sum_{x^{c}\leq p<x}\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ p|p^{\prime}-1\end{subarray}}1=\sum_{x^{c}\leq p<x}\sum_{\begin{subarray}{c}\ell p+1\leq x\\ \ell p+1\in\mathcal{P}\\ 2|\ell\end{subarray}}1\leq\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\sum_{\begin{subarray}{c}\ell p\leq x\\ \ell p+1\in\mathcal{P}\end{subarray}}1.$
@@
where $c$ is a fixed number satisfying $0.75<c<1$. Let $\mathcal{P}=\mathbb{P}\setminus\{2\}$ and define the sieve function $S(\mathcal{A},\mathcal{P},z)$ to be
$S(\mathcal{A},\mathcal{P},z):=\big\{a\in\mathcal{A}:\big(a,P(z)\big)=1\big\},$

### v2

@@ Abstract.
We denote by $P^{+}(n)$ the largest prime factor of the integer $n$. In 1935, Erdős studied the quantity $T_{c}(x)$ defined by
$T_{c}(x)=\big|\big\{p\leq x:P^{+}(p-1)\geq p^{c}\big\}\big|,$
and he proved
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\rightarrow 0,\quad\text{as~}c\rightarrow 1.$
Recently, Ding gave a quantitative form of Erdős’ result, showing that
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\leq 8\big(c^{-1}-1\big).$
holds for $8/9<c<1$. In this paper, we improve Ding’s upper bound to
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\leq-\frac{7}{2}\log c$
@@ 1. Introduction
$T_{c}(x):=\big|\big\{p\leq x:P^{+}(p-1)\geq p^{c}\big\}\big|,$
where $0<c<1$.
For small values of $c\leq 1/2$, here we just list several results on $T_{c}(x)$. In 2015, Luca, Menares, and Pizarro-Madariaga [25] showed that
$\liminf_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\geq 1-c$ (1.1)
holds for $1/4\leq c\leq 1/2$. Later, Chen and Chen [8] extended the range of $c$ to $(0,1/2)$ for (1.1). In 2018, Feng and Wu [14] improved the lower bound in (1.1) for $0<c<0.3517\ldots$ by showing
$\liminf_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\geq 1-4\int_{1/c-1}^{1/c}\frac{\rho(t)}{t}dt,$
where $\rho(u)$ is the Dickman function. Later, the above lower bound was further improved by Liu, Wu and Xi [24] to
$\liminf_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\geq 1-4\rho(1/c)$
provided $0<c<0.3734\ldots$. Finally, we refer the readers to e.g. Banks and Shparlinski [4], Wu [31, 32] for more results related to this topic.
Now we turn our attention to the case when $1/2<c\leq 1$ is large. It seems this case is more difficult and more attractive, considering its close connection with the twin prime conjecture and Fermat’s last theorem as we have mentioned above. On the one hand, one may want to seek values of $c$ as large as possible such that
$\liminf_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}>0.$ (1.2)
In 1969, Goldfeld [17] proved that (1.2) holds for some $c>1/2$ by employing the Bombieri-Vinogradov theorem as well as the Brun-Titchmarsh inequality, and he also stated that his arguments is also applicable for $c\approx 7/12$. Since then, there has been a number of improvements on the lower bound of values of $c$ in (1.2), see, e.g. Motohashi [27], Hooley [20, 21], Deshouillers and Iwaniec [10], Fouvry [15], Baker and Harman [2]. The best record of $c$ up to now is $0.677$, obtained by Baker and Harman [3].
On the other hand, one may want to have a non-trivial upper bound for $T_{c}(x)$ for some fixed large $c$, which is the main topic we shall study in this paper. In fact, we even conditionally know the asymptotic formula for $T_{c}(x)$:
$\lim_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}=1-\rho\left(\frac{1}{c}\right)$ (1.3)
holds for $0<c<1$ under the assumption of the Elliott–Halberstam conjecture, see e.g. the works of Pomerance [28], Granville [18], Wang [29] and Wu [33]. Motivated by this, in 2023, Ding [11] proved unconditionally that
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}<1/2$ (1.4)
for some absolute constant $c<1$. The proof of (1.4) by Ding is based on the following result of Brun-Titchmarsh type due to Wu [33, Lemma 2.2].
@@
For $(a,m)=1$, let $\pi(x;m,a)$ denotes the number of primes $p\leq x$ such that $p\equiv a\pmod{m}$.There exist two functions $K_{2}(\theta)>K_{1}(\theta)>0$, defined on the interval $(0,17/32)$ such that for each fixed $A>0$, and sufficiently large $Q=x^{\theta}$, the inequalities
@@
The constant $c$ in (1.4) could further be specified by explicit values of $K_{1}(\theta)$ in Proposition 1.1, for example, one has $K_{1}(\theta)\geq 0.16$ for $1/2\leq\theta\leq 13/25$ [2, Theorem 1]. Using the method of Ding [11] as well as the explicit values of $K_{1}(\theta)$, Xinyue Zang (private communication) obtained that
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\leq\frac{0.496875}{c}<\frac{1}{2},\quad\text{for~}0.993375<c<1.$ (1.5)
However, there are earlier results related to (1.4) and (1.5). Actually, as indicated by the proof of a former result of Erdős [13, from line -6, page 212 to line 4, page 213], as early as 1935, one could already conclude that
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\rightarrow 0,\quad\text{as}~c\rightarrow 1,$ (1.6)
which just corresponds to the asymptotic (1.3) since $1-\rho(1/c)\to 0$ as $c\to 1$. Clearly, (1.4) is now a simple corollary of (1.6). Very recently, Ding [12] obtained a quantitative form of Erdős’ result (1.6):
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\leq 8\big(c^{-1}-1\big)$ (1.7)
@@
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}<\frac{1}{2}$ (1.8)
@@
[19, Theorem 5.7] Let $g$ be a natural number, and let $a_{i},b_{i}~(i=1,2,\cdot\cdot\cdot,g)$ be integers satisfying
@@
In this article, we shall give a further improvement of (1.7) with two new ingredients: the first is the employment of linear sieve to the prime variable sequence, instead of integer variable polynomial combining with the two dimensional sieve; the second one is that when dealing with the error term coming from linear sieve, we apply a theorem of Bombieri-Friedlander-Iwaniec type with level of distribution $x^{4/7-\varepsilon}$ instead of the classical level $x^{1/2-\varepsilon}$.
@@
$\limsup_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}\leq-\frac{7}{2}\log c.$
@@
Remark 2.
Recall that $\rho(u)=1$ for $0\leq u\leq 1$ and $\rho(u)=1-\log u$ for $1\leq u\leq 2$ from the recursion formula (see e.g. [26, Eq. (7.6)]) on Dickman’s function. Thus, by (1.3) we have conditionally for $1/2<c<1$
$\lim_{x\rightarrow\infty}\frac{T_{c}(x)}{\pi(x)}=1-\rho\left(\frac{1}{c}\right)=-\log c$
under the assumption of Elliott–Halberstam conjecture, which seems far out of reach at present. Our unconditionally bound $\frac{7}{2}\log c$ in Theorem 1.1 is 3.5 times that of the expected asymptotic density.
The following corollary of Theorem 1.1 improves (1.4), (1.5) and (1.8) considerably.
@@
The numerical value of $e^{-\frac{1}{7}}$ is $\approx 0.86687$. According to the conditional asymptotic formula (1.3), one should expect $\limsup_{x\rightarrow\infty}T_{c}(x)/\pi(x)<1/2$ for any $c>e^{-\frac{1}{2}}\approx 0.60653$. Corollary 1.1 makes some further progress toward this direction.
@@ 2. Fundamental lemmas
Let $\mu(n)$ be the Möbius function. Let $\mathcal{A}$ be a finite sequence of positive integers and let $\mathcal{P}$ be a subset of primes. For any $z\geq 2$, we set
@@
The first lemma is due to Iwaniec, see [22] or [23].
Lemma 2.1.
Let $D\geqslant z\geqslant 2$ and $L>1$. Put $\lambda_{1}^{\pm}=1$ and for all squarefree numbers $d=p_{1}\cdots p_{r},\,p_{1}>\cdots>p_{r},\,r\geq 1$, define the upper and lower bound sieve weights of level $D$ respectively
$\ \qquad\lambda_{d}^{+}=\left\{\begin{array}[]{ll}(-1)^{r}&\ \ {\rm if}\ p_{{}_{1}}\cdots p_{{}_{2l}}p_{{}_{2l-1}}^{3}<D\ {\rm for\ all\ }0\leq l\leq(r-1)/2,\\ \vskip 5.69054pt\cr 0&\ \ {\rm otherwise},\end{array}\right$.
and
$\lambda_{d}^{-}=\left\{\begin{array}[]{ll}(-1)^{r}&\ \ {\rm if}\ p_{{}_{1}}\cdots p_{{}_{2l-1}}p_{{}_{2l}}^{3}<D\ {\rm for\ all\ }0\leq l\leq r/2,\\ \vskip 5.69054pt\cr 0&\ \ {\rm otherwise}.\end{array}\right$.
Then we have
$\lambda_{d}^{-}*\mathbf{1}\leqslant\mu*\mathbf{1}\leqslant\lambda_{d}^{+}*\mathbf{1}$
and
$\sum_{d\mid P(z)}\lambda^{+}_{d}\frac{\omega(d)}{d}$ $\leqslant\prod_{p\leqslant z\atop p\in\mathcal{P}}\left(1-\frac{\omega(p)}{p}\right)\left\{F(s)+O\left(\frac{e^{\sqrt{L}-s}}{(\log D)^{1/3}}\right)\right\},$
$\sum_{d\mid P(z)}\lambda^{-}_{d}\frac{\omega(d)}{d}$ $\geqslant\prod_{p\leqslant z\atop p\in\mathcal{P}}\left(1-\frac{\omega(p)}{p}\right)\left\{f(s)-O\left(\frac{e^{\sqrt{L}-s}}{(\log D)^{1/3}}\right)\right\}$
uniformly for all multiplicative functions $w$ satisfying
${\rm{(i)}}\quad 0<\omega(p)<p\quad(p\in\mathcal{P}),$
${\rm{(ii)}}\prod_{\begin{subarray}{c}u<p\leqslant v\\[2.84526pt] p\in\mathcal{P}\end{subarray}}\left(1-\frac{\omega(p)}{p}\right)^{-1}\leqslant\frac{\log v}{\log u}\left(1+\frac{L}{\log u}\right)\quad(2\leqslant u\leqslant v\leqslant z).$
where $s=\log D/\log z$ and $F(s),f(s)$ are defined by the continuous solutions to the system
$\left\{\begin{array}[]{ll}sF(s)=2\mathrm{e}^{\gamma}&\ \ 1\leq s\leq 2,\\ \vskip 5.69054pt\cr sf(s)=0&\ \ 0<s\leq 2,\\ \vskip 5.69054pt\cr(sF(s))^{\prime}=f(s-1)&\ \ s>2,\\ \vskip 5.69054pt\cr(sf(s))^{\prime}=F(s-1)&\ \ s>2.\end{array}\right$.
Here $\gamma$ is the Euler constant.
The second lemma is a convolution generalization theorem of Bombieri-Friedlander-Iwaniec [6] with well-factorable weights. An arithmetic function $\lambda(q)$ is called well factorable of level $Q$ if for any $Q=Q_{1}Q_{2}$, $Q_{1},Q_{2}\geq 1$, there exist two functions $\lambda_{1}$ and $\lambda_{2}$ supported in $1,Q_{1}]$ and $1,Q_{2}]$ respectively such that
@@
Next for given integers $\ell,a,q$ with $(a,q)=1$, we define the convolution sum
$\pi(y;\ell,a,q):=\sum_{\begin{subarray}{c}\ell p\leq y\\ \ell p\equiv a\!\!\!\!\!\pmod{q}\end{subarray}}1,$
where the symbols $p$ will always be primes. Then we have
@@
$\sum_{\begin{subarray}{c}q\leq Q\\ (a,q)=1\end{subarray}}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leq\ell\leq L_{2}\\ (\ell,q)=1,~2|\ell\end{subarray}}\bigg(\pi(x;\ell,a,q)-\frac{\pi\big(x/\ell\big)}{\varphi(q)}\bigg)\ll\frac{x}{(\log x)^{A}}$
holds for $Q\leq x^{4/7-\varepsilon}$ and $1\leq L_{1}\leq L_{2}\leq x^{1-\varepsilon}$, where the implied constants depend only on $a,A$ and $\varepsilon$, $\varphi(n)$ is the Euler totient function.
@@
In fact, the original statement of Wang’s proposition [30, Proposition 3.2] is slightly different from Lemma 2.2. For our applications, we add the additional restriction $2|\ell$ here in Lemma 2.2. The proof is almost the same as Wang [30, Proposition 3.2]. After applying Heath-Brown’s identity, we shall consider the sum
@@
$\sum_{(a,\,q)=1}\lambda(q)\sum_{\begin{subarray}{c}L_{1}\leqslant\ell\leqslant L_{2}\\ (\ell,\,q)=1\end{subarray}}f(\ell)\bigg(\pi(x;\,\ell,a,q)-\frac{\pi(x/\ell)}{\varphi(q)}\bigg)\ll_{a,A,\varepsilon}\frac{x}{(\log x)^{A}}$
@@
The last lemma is another conjecture of Chen and Chen [8] which was later confirmed by Wu [33, Theorem 2].
@@
$T^{\prime}_{c}(x)=\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ P^{+}(p^{\prime}-1)\geq x^{c}\end{subarray}}1=\sum_{x^{c}\leq p<x}\sum_{\begin{subarray}{c}p^{\prime}\leq x\\ p|p^{\prime}-1\end{subarray}}1=\sum_{x^{c}\leq p<x}\ \sum_{\begin{subarray}{c}\ell p+1\leq x\\ \ell p+1\,\textrm{is\, prime}\\ 2|\ell\end{subarray}}1\leq\sum_{\begin{subarray}{c}\ell\leq x^{1-c}\\ 2|\ell\end{subarray}}\sum_{\begin{subarray}{c}\ell p\leq x\\ \ell p+1\,\textrm{is\, prime}\end{subarray}}1.$
@@
for some fixed $c$, $0.75<c<1$. Let $\mathcal{P}=\{p:\ p\neq 2\}$ and define the sieve function $S(\mathcal{A},\mathcal{P},z)$ by
$S(\mathcal{A},\mathcal{P},z)=\big\{a\in\mathcal{A}:\big(a,P(z)\big)=1\big\},$